OpenStudy (anonymous):

Prove: tanx/(secx+1)=(secx-1)/tanx tanx/secx+1=?

Parth (parthkohli):

Is this the equation?\[\dfrac{\tan(x)}{\sec(x) }+1 = \dfrac{\sec(x) - 1}{\tan(x)}\]

OpenStudy (anonymous):

i'll fix it

OpenStudy (anonymous):

itit should be showing up properly now

Parth (parthkohli):

Though you may keep in mind that \(\tan(x) = \dfrac{\sin(x)}{\cos(x)}\) and \(\sec(x) = \dfrac{1}{\cos(x)}\)

OpenStudy (whpalmer4):

I think it is \[\frac{\tan x}{\sec x + 1} = \frac{\sec x - 1}{\tan x}\] I would cross multiply to get the ball rolling...

Parth (parthkohli):

Ah, a very good idea.

OpenStudy (anonymous):

you can't use both sides to prove it

OpenStudy (whpalmer4):

Sure you can. You manipulate the equation in any legitimate fashion (in other words, do the same thing to both sides) until you end up with an equation which is clearly true.

Parth (parthkohli):

You can. Assume that it is true, and then use them. If you get to a point where everything is obvious, then it is true.

OpenStudy (whpalmer4):

Similarly, if you end up at a point where everything is obviously not true, then you've proven that it is not true.

OpenStudy (anonymous):

ahh

OpenStudy (whpalmer4):

My rough sketch of how to prove this one, without remembering any obscure identities: cross-multiply (notice you get a difference of two squares for one side) rewrite sec as 1/cos rewrite tan as sin/cos multiply through by what hopefully is an obvious choice use the most basic trig identity: \(\sin^2 x+ \cos^2 x = 1\)

Parth (parthkohli):

You can even use the identity \(\tan^2(x) = 1 - \sec^2(x)\) directly.

OpenStudy (whpalmer4):

I'd have to prove that one to myself before using :-) Never did like memorizing trig identities!

OpenStudy (anonymous):

^ I know that feeling

Parth (parthkohli):

Me too :-P

OpenStudy (anonymous):

Alright lemme see if I can get to the answer lol

OpenStudy (jhannybean):

Me 3.

OpenStudy (anonymous):

so i somehow got to sin(x)/sin^2x+cos^2x

OpenStudy (jhannybean):

I keep my calculus book close to my heart...only because it's got all the trig identities and stuff in back, lolol.

Parth (parthkohli):

LOL

OpenStudy (whpalmer4):

well, you need a bit more than that, there's no = sign!

OpenStudy (whpalmer4):

The URL for a trig identity page on my iPad takes up much less space than my old calc book :-)

Parth (parthkohli):

iPad? Richie rich alert.

OpenStudy (jhannybean):

I can use my calc book as a weapon. -_- it's gigantic.

OpenStudy (anonymous):

man im so confuffled

OpenStudy (whpalmer4):

Can you show us what you have?

Parth (parthkohli):

@isuckatmath9999 Hint: 2 + 2 = 4.

Parth (parthkohli):

Use that hint and you can get your answer.

OpenStudy (anonymous):

How is that a hint?

Parth (parthkohli):

You can use 2 + 2 = 4, then derive arithmetic. Then derive algebra from it. Then derive trigonometry from it, and solve your question :3

Parth (parthkohli):

Hehe. Just cross-multiply, eh?

OpenStudy (whpalmer4):

Here, I'll show you the first two steps: 1) write "2+2" 2) write "=4" :-)

OpenStudy (anonymous):

I have some gibberish of this tan(x)/(secx-1) =(sin(x)/cos(x))*(cos(x)/1) =(sin(x)/1) =sin(x)/sin^2x+cos^2x

Parth (parthkohli):

\[\tan^2(x) = (\sec(x) - 1)(\sec(x) + 1)\]Simplify the right hand side using a well-known identity

Parth (parthkohli):

You shouldn't write tan(x) = sin(x)/cos(x) right from the first step. That'd make things harder for you.

OpenStudy (anonymous):

>.>

Parth (parthkohli):

As I said, simplify \(\tan^2(x) = (\sec(x) - 1)(\sec(x) + 1)\)

Parth (parthkohli):

The right-hand side, specifically.

OpenStudy (anonymous):

theres no tan^2x on the right side is there? what?

Parth (parthkohli):

Uh no. I'm telling you to simplify \((\sec(x) - 1)(\sec(x) + 1)\)

OpenStudy (whpalmer4):

Seriously, after cross multiplying, you should have \[\tan^2 x = (\sec x -1)(\sec x + 1) = \sec^2 x - 1\]\[\tan^2 x = \frac{1}{\cos^2 x} - 1\]Now plug in \[\tan x = \frac{\sin x}{\cos x}\]\[\frac{\sin^2x}{\cos^2x} =\frac{1}{\cos^2 x} - 1 \]and it should be easy from there

OpenStudy (whpalmer4):

Just pretend you're simplifying \[\frac{s^2}{c^2} = \frac{1}{c^2}-1\] and replace s and c by sin and cos when you're done.

OpenStudy (anonymous):

*twitch*

OpenStudy (whpalmer4):

Or don't. If you don't, how about multiplying everything by \(\cos^2 x\) and see what happens?

OpenStudy (anonymous):

this isn't helping me at all. =/

OpenStudy (whpalmer4):

Do you see how I got to \[\frac{\sin^2 x}{\cos^2 x} = \frac{1}{\cos^2 x} - 1\]?

OpenStudy (whpalmer4):

I'll start again at the top, and you tell me after each step if you've got it: \[\frac{\tan x}{\sec x + 1} = \frac{\sec x - 1}{\tan x}\]We cross-multiply: \[\tan x * \tan x = (\sec x -1)(\sec x + 1) \]\[\tan^2 x = \sec^2x-\sec x + \sec x -1 = \sec^2 x - 1\]\[\tan^2 x = \sec^2 x -1\] Okay so far?

OpenStudy (whpalmer4):

Substitute \[\sec x = \frac{1}{\cos x}\]\[\tan^2 x = \frac{1}{\cos^2 x} - 1\] Substitute \[\tan x = \frac{\sin x}{\cos x}\]\[\frac{\sin^2 x}{\cos^2 x} = \frac{1}{\cos^2 x} -1\]Multiply everything by \(\cos^2x\) \[\frac{\sin^2x}{\cancel{\cos^2x}}*\cancel{\cos^2x} = \frac{1}{\cancel{\cos^2x}}*\cancel{\cos^2 x} - \cos^2 x\]\[\sin^2 x= 1 -\cos^2x\]Add \(\cos^2x\) to both sides \[\sin^2x + \cos^2 x = 1\checkmark\]